# Light-induced localized vortices in multicomponent Bose–Einstein condensates

### v1

@@ III.2 Full treatment
Figure 4: Same as Fig. 3 for $\nu=1$ and $a=0.02$.
@@
We will now turn to the solutions of the full coupled GPEs (10), including the atom–atom interactions, and study the structure of the lowest-energy states of the dark-state manifold. To obtain the numerical solutions, we employed the method of imaginary-time evolution: one makes a change $t\to-\mathrm{i}\tau$ and propagates the equations in time $\tau$ until convergence is reached, starting from a certain trial wave function. The algorithm will then converge to the lowest-energy state provided it has nonzero overlap with the trial wave function. To enable convergence to the states of the dark-state manifold, we used the solution of Eq. (13) as the trial function for solving coupled GPEs (10). In turn, the trial wave functions required to solve Eq. (13) were constructed by solving this equation using the Thomas–Fermi approximation and multiplying the solution by a phase factor $\mathrm{e}^{\mathrm{i}l\varphi}$ with a chosen value of $l$. Since states of the form (18) with different $l$ are orthogonal, the trial wave function of this form converges to a state with the same value of $l$. This allowed us to obtain stationary solutions of different vorticities. The ground state can then be found by ordering the states based on the value of the chemical potential $\mu$ (or based on the energy per particle $E$). We refer the reader to the Appendix for additional numerical details.
@@
We start by setting $\nu=1$ and $a=0.5$. Such a value of $a$ corresponds to an intermediate regime whereby the term $\rho A_{\varphi}$ is not small (compared to unity) and cannot be approximated by $-\nu$, as one can see in the lower panel of Fig. 3(a). Numerical calculations show that ground state is the one corresponding to an $l=0$ solution of the dark-state GPE (13). As noted above, this corresponds to $\Phi_{1}$ being vortex-free and $\Phi_{2}$ being in a vortex state. Such a solution is shown in Fig. 3(b). It is apparent that the first component fills the core of the second component, a situation known to stabilize vortices with vorticities higher than $l=1$ [35]. Notably, which of the components acquires vorticity is determined by the inequivalent coupling fields $\Omega_{1}$ and $\Omega_{2}$ rather than the intrinsic properties of the two states (the scattering lengths).
@@
Two lowest excited states are shown in Figs. 3(c)–(d). Their energies per particle are 96.7 and 103.4 units, respectively, showing that the term $\rho A_{\varphi}$ has lifted the degeneracy. Notably, the “interaction potentials” $g_{ij}|\Phi_{k}|^{2}$ do not depend on the vorticities of the components and, therefore, do not lift the degeneracy. In the first excited state, the first component is in a vortex state, rotating in the direction opposite to the direction of the vector field $\boldsymbol{A}$, while the second component is vortex-free. The vortex on the first component mostly occupies the space at the center of the second component. Such an approximate phase separation is in line with the criterion $g_{12}^{2}>g_{11}g_{22}$ [36, 37] which, however, does not take into account the kinetic energy [38], which is especially important in the present setting. In the second excited state, vorticities of the two components are given by $l_{1}=1$ and $l_{2}=2$, while the spatial profiles are almost the same as those of the respective components of the first excited state.
@@
Let us now study the regime where the distance $a$ is much smaller that the system radius $R$, specifically $a=0.02$. In this case the condition $\rho\gg a$ holds in most of the space, and $\rho A_{\varphi}$ tends to $-\nu$ already in the vicinity of the origin [see Fig. 4(a), lower panel]. The ground state now corresponds to the $l=-\nu=-1$ solution of the dark-state GPE (13). This means that $\Phi_{1}$ is a vortex state, and it occupies a small region of space near the origin [see Fig. 4(b)]. Specifically, the density of the first component is concentrated in a region of the order of the healing length $\xi_{2}$ of the second component. Assuming all atoms are in the second component, we have (in dimensionless units) $\xi_{2}=\sqrt{S/g_{22}}=0.086$, where we approximately took $S=\pi$ for the occupied area. The localized nature of the vortex results from the $\frac{1}{\sqrt{1+|\zeta|^{2}}}$ term in the expression for $\Phi_{1}$ (17), which effectively cuts the vortex off when $\rho\gg a$. Another consequence is the small resulting population of the first component, which in the present case is $\eta_{1}=0.25\%$. Thus, decreasing the $a$ parameter localizes the vortex more tightly, but also reduces the population of the first component. The second component occupies the space surrounding the first component and has zero vorticity.
@@
The first excited state, on the other hand, has the typical structure whereby the non-rotating component fills the core of the rotating one [see Fig. 4(c)]. The almost-degenerate partner of this state is shown in Fig. 4(d). These two states correspond, respectively, to the $l=0$ and $l=-2$ solutions of the dark-state GPE (13). The degeneracy results from the fact that the kinetic energy is the same, whether the cloud does not rotate ($l=0$) or rotates in the direction opposite to the $\boldsymbol{A}$ field but with twice the speed ($l=-2$).
@@
Localized vortices with vorticities higher than one can be analogously created by using beams with $\nu>1$. The results for the case $\nu=2$ are presented in Fig. 5. Choosing $a=0.1$ already leads to the potential $\rho A_{\varphi}$ tending to $-\nu$ in most of space [see Fig. 5(a)]. This leads to a ground state with the first component having vorticity $l_{1}=-2$ and the second being vortex-free: $l_{2}=l_{1}+\nu=0$. The vortex has a density maximum farther from origin compared to the $l_{1}=-1$ vortex in Fig. 4(b) due, in part, to the former having the form of $J_{2}$ and thus rising less sharply compared to $J_{1}$. The degree of localization can be increased by reducing $a$, but the number of atoms in the vortex component will then decrease. The lowest excited states [see Figs. 5(c)–(d)] also contain localized vortices in the first component.
@@
Figure 6: Wave functions of components 1 and 2 after completing one circular sweep of the laser beams around the origin. The tangential beam movement speed is (a) $v_{1}=150\,\mathrm{\mu m/s}$, (b) $v_{2}=300\,\mathrm{\mu m/s}$, and (c) $v_{3}=800\,\mathrm{\mu m/s}$. In all three cases the relative population of the first component is $\eta_{1}\approx 0.0025$.
@@ IV Moving the vortex
To demonstrate the stability of the localized lowest-energy vortex states and show the high level of available control, we consider changing the position of the vortex by moving the laser beams. A similar protocol has been considered in [39], however, that work considered moving a vortex created by a phase-imprinting technique. Our approach is more robust because the laser beams create a pinned vortex which cannot break free.
@@
We focus on the regime $\nu=1$, $a=0.02$ studied in Fig. 4 but this time consider an off-axis vortex created by centering the laser beams at $(x,y)=(0.5,0)$. Once the state is prepared, the transverse profile of the laser beams $\Omega_{1}$ and $\Omega_{2}$ start moving on a circular path around the origin at a constant tangential speed $v$. Figure 6 shows the wave functions obtained after completing one full circle. In Fig. 6(a) corresponding to $v_{1}=150\,\mathrm{\mu m/s}$ (time to complete the circle is $t=0.314\,\mathrm{s}$), it is apparent that the vortex of the first component has retained its structure, and the density of the second component has not been strongly distorted. The value of $E$ has increased by only $3\%$ of the energy (per particle) gap between the ground state and the first excited one. Here the velocity is chosen to be lower than the speed of sound in the second component, which equals to $s=400\,\mathrm{\mu m/s}$. Moving the vortex at twice the speed, $v_{2}=2v_{1}$, results in setting the whole of the second component in circular motion. The distortion of the density is apparent in Fig. 6(b).
@@
Finally, the case of “supersonic” movement is studied in Fig. 6(c) corresponding to the movement speed $v_{3}=2s$. The density profiles of both components are strongly distorted and the energy is increased 1.5 times. Nevertheless, the first component retains its signature phase profile, and no additional vortices have been created in the second component in the process.
@@ V Conclusions
In summary, we have studied the interaction of a two-component BEC mixture with the light fields in a $\Lambda$-type configuration and investigated the stationary states of the dark-state manifold. The angular momentum of $\nu\hbar$ per photon carried by one of the two fields leads to either one or both components being in a vortex state, with their vorticities differing by $\nu$ units. In the regime $a\ll R$, the structure of the stationary states is dictated by the vector potential term. In this limit, $\boldsymbol{A}$ tends to $-\frac{\nu}{\rho}\boldsymbol{e}_{\varphi}$ and its contribution to the velocity field (and hence to the kinetic energy) can be canceled if the dark state is described by a wave function $f(\rho)\mathrm{e}^{-\mathrm{i}\nu\varphi}$. We have demonstrated that the lowest-energy state of the dark-state manifold indeed has this form. In this case, the first component (the one interacting with the LG beam) contains a vortex of vorticity $\nu$, while the second component is vortex-free. The core of the former vortex coincides with the center of the beams, while the density profile of the vortex demonstrates a strong degree of localization as the density falls off as $[1+(\rho/a)^{2}]^{-1/2}$ away from the vortex core. Such a vortex can be moved around by moving the laser beams. Provided the movement speed is less than approximately half the speed of sound in the condensate, the shape of the vortex retains its structure during the movement, and the density of the second component does not get distorted.
@@ Acknowledgements.
Y. B. would like to acknowledge the outstanding hospitality received during his internship at the Okinawa Institute of Science and Technology Graduate University, Quantum Systems Unit. This project was supported by the LRC grant No. S-LJB-24-2 and the JSPS Bilateral Program No. JPJSBP120244202. This work has also received funding from COST Action POLYTOPO CA23134, supported by COST (European Cooperation in Science and Technology).
@@
The wave functions are normalized to unity: $\sum_{i=1}^{3}\int\mathrm{d}\boldsymbol{r}\,|\Phi_{i}|^{2}=1$.
@@
$\begin{split}\boldsymbol{A}&=-\frac{\nu}{a}\frac{\left(\frac{\rho}{a}\right)^{2\nu-1}}{1+\left(\frac{\rho}{a}\right)^{2\nu}}\boldsymbol{e}_{\phi},\\ U&=2\left(\frac{\nu}{a}\right)^{2}\frac{\left(\frac{\rho}{a}\right)^{2\nu-2}}{[1+\left(\frac{\rho}{a}\right)^{2\nu}]^{2}}.\end{split}$ (A4)

### v2

@@ III.2 Full treatment
Figure 3: The potentials $U$ and $\rho A_{\varphi}$, as well as the solutions of Eq. (10) for $\nu=1$ and $a=0.5$. (a) Potentials $U$ and $\rho A_{\varphi}$. (b), (c), and (d) show, respectively, three lowest-energy solutions of Eq. (10); wave function normalization is $\sum_{i=1}^{3}\int_{S}\,|\psi_{i}|^{2}\,\mathrm{d}S=1$. The black lines in the upper and lower panels show, respectively, the radial cuts of the densities $|\psi_{1}|^{2}$ and $|\psi_{2}|^{2}$ of the first and the second component. The gray lines show the total density. The values of the density of the third component are vanishingly small in all cases (calculation yields values no larger than $10^{-7}$) and are therefore not shown. For the same reason, the relative occupation of the second component can be taken to be $\eta_{2}=1-\eta_{1}$. The vorticities $l_{i}$ of the two components are displayed together with the insets showing the two-dimensional phase profiles in the xyxy-plane. The value of the phase is color-coded as follows: $\text{dark blue}=-\pi$, $\text{white}=0$, $\text{dark red}=\pi$.
Figure 4: Same as Fig. 3 for $\nu=1$ and $a=0.02$. The total densities are not explicitly shown since they would appear to largely coincide with $|\psi_{2}|^{2}$ due to the highly imbalanced occupation of the two modes.
@@
We will now turn to the solutions of the full coupled GPEs (10), including the atom–atom interactions, and study the structure of the lowest-energy states of the dark-state manifold. To obtain the numerical solutions, we employed the method of imaginary-time evolution: one makes a change $t\to-\mathrm{i}\tau$ and propagates the equations in time $\tau$ until convergence is reached, starting from a certain trial wave function. The algorithm will then converge to the lowest-energy state provided it has nonzero overlap with the trial wave function. To enable convergence to the states of the dark-state manifold, we used the solution of Eq. (13) as the trial function for solving coupled GPEs (10). In turn, the trial wave functions required to solve Eq. (13) were constructed by solving this equation using the Thomas–Fermi approximation and multiplying the solution by a phase factor $\mathrm{e}^{\mathrm{i}l\varphi}$ with a chosen value of $l$. Since states of the form (18) with different $l$ are orthogonal, the trial wave function of this form converges to a state with the same value of $l$. This allowed us to obtain stationary solutions of different vorticities. The ground state can then be found by ordering the states based on the value of the energy per particle $E$. The population of each component was calculated from the obtained solutions rather than being enforced. We refer the reader to the Appendix for additional numerical details.
@@
We start by setting $\nu=1$ and $a=0.5$. Such a value of $a$ corresponds to an intermediate regime whereby the term $\rho A_{\varphi}$ is not small (compared to unity) and cannot be approximated by $\nu$, as one can see in the lower panel of Fig. 3(a). Numerical calculations show that ground state is the one corresponding to an $l=0$ solution of the dark-state GPE (13). As noted above, this corresponds to $\Phi_{1}$ being vortex-free and $\Phi_{2}$ being in a vortex state. Such a solution is shown in Fig. 3(b). It is apparent that the first component fills the core of the second component, a situation known to stabilize vortices with vorticities higher than $l=1$ [35, 36]. Notably, which of the components acquires vorticity is determined by the inequivalent coupling fields $\Omega_{1}$ and $\Omega_{2}$ rather than the intrinsic properties of the two states (the scattering lengths). The total density of the BEC, shown in gray in Fig. 3(b), indicates that the condensate as a whole does not have a pronounced density dip at the origin, meaning that the core of the vortex is “completely” filled.
@@
Two lowest excited states are shown in Figs. 3(c)–(d). Their energies per particle are 96.7 and 103.4 units, respectively, showing that the term $\rho A_{\varphi}$ has lifted the degeneracy. Notably, the “interaction potentials” $g_{ij}|\Phi_{k}|^{2}$ do not depend on the vorticities of the components and, therefore, do not lift the degeneracy. In the first excited state, the first component is in a vortex state, rotating in the direction opposite to the direction of the vector field $\boldsymbol{A}$, while the second component is vortex-free. The vortex on the first component mostly occupies the space at the center of the second component. Such an approximate phase separation is in line with the criterion $g_{12}^{2}>g_{11}g_{22}$ [37, 38] which, however, does not take into account the kinetic energy [39], which is especially important in the present setting. In the second excited state, vorticities of the two components are given by $l_{1}=1$ and $l_{2}=2$, while the spatial profiles are almost the same as those of the respective components of the first excited state. In both cases, the components contribute to a nearly uniform total density of the cloud in the bulk.
@@
Let us now study the regime where the distance $a$ is much smaller that the system radius $R$, specifically $a=0.02$. In this case the condition $\rho\gg a$ holds in most of the space, and $\rho A_{\varphi}$ tends to $\nu$ already in the vicinity of the origin [see Fig. 4(a), lower panel]. The ground state now corresponds to the $l=-\nu=-1$ solution of the dark-state GPE (13). This means that $\Phi_{1}$ is a vortex state, and it occupies a small region of space near the origin [see Fig. 4(b)]. Specifically, the density of the first component is concentrated in a region of the order of the healing length $\xi_{2}$ of the second component. Assuming all atoms are in the second component, we have (in dimensionless units) $\xi_{2}=\sqrt{S/g_{22}}=0.086$, where we approximately took $S=\pi$ for the occupied area. The localized nature of the vortex results from the $\frac{1}{\sqrt{1+|\zeta|^{2}}}$ term in the expression for $\Phi_{1}$ (17), which effectively cuts the vortex off when $\rho\gg a$. Another consequence is the small resulting population of the first component, which in the present case is $\eta_{1}=0.25\%$. Thus, decreasing the $a$ parameter localizes the vortex more tightly, but also reduces the population of the first component. The second component occupies the space surrounding the first component and has zero vorticity.
@@
The first excited state, on the other hand, has the typical structure whereby the non-rotating component fills the core of the rotating one [see Fig. 4(c)]. The almost-degenerate partner of this state is shown in Fig. 4(d). These two states correspond, respectively, to the $l=0$ and l=2l=-2 solutions of the dark-state GPE (13). The degeneracy results from the fact that the kinetic energy is the same, whether the cloud does not rotate ($l=0$) or rotates in the direction opposite to the $\boldsymbol{A}$ field but with twice the speed (l=2l=-2).
@@
Localized vortices with vorticities higher than one can be analogously created by using beams with $\nu>1$. The results for the case $\nu=2$ are presented in Fig. 5. Choosing $a=0.1$ already leads to the potential $\rho A_{\varphi}$ tending to $\nu$ in most of space [see Fig. 5(a)]. This leads to a ground state with the first component having vorticity $l_{1}=-2$ and the second being vortex-free: $l_{2}=l_{1}+\nu=0$. The vortex has a density maximum farther from origin compared to the $l_{1}=-1$ vortex in Fig. 4(b) due, in part, to the former having the form of $J_{2}$ and thus rising less sharply compared to $J_{1}$. The degree of localization can be increased by reducing $a$, but the number of atoms in the vortex component will then decrease. The lowest excited states [see Figs. 5(c)–(d)] also contain localized vortices in the first component.
It is worth noting that the density profiles of both components remain essentially unchanged as $g_{12}$ is reduced into the miscible regime, $g_{12}^{2}<g_{11}g_{22}$, provided the coupling constants $g_{ij}$ are of similar magnitude. In particular, the densities shown in Figs. 4(b) and 5(b) are virtually unaffected. This invariance is expected, as the localized-vortex configuration is sustained by the external laser beams rather than interparticle interactions, as demonstrated in Section III A, where interactions are entirely neglected.
@@
Figure 6: Wave functions of components 1 and 2 after completing one circular sweep of the laser beams around the origin. The tangential beam movement speed is (a) $v_{1}=3v_{\rm s}/8=150\,\mathrm{\mu m/s}$, (b) $v_{2}=3v_{\rm s}/4=300\,\mathrm{\mu m/s}$, and (c) $v_{3}=2v_{\rm s}=800\,\mathrm{\mu m/s}$. In all three cases the relative population of the first component is $\eta_{1}\approx 0.0025$, while the third component is essentially unpopulated.
@@ IV Moving the vortex
To demonstrate the stability of the localized lowest-energy vortex states and show the high level of available control, we consider changing the position of the vortex by moving the laser beams. A similar protocol has been considered in Ref. [40], however, that work considered moving a vortex created by a phase-imprinting technique. Our approach is more robust because the laser beams create a pinned vortex which cannot break free.
@@
We focus on the regime $\nu=1$, $a=0.02$ studied in Fig. 4 but this time consider an off-axis vortex created by centering the laser beams at $(x,y)=(0.5,0)$. Once the state is prepared, the transverse profile of the laser beams $\Omega_{1}$ and $\Omega_{2}$ start moving counterclockwise on a circular path around the origin at a constant tangential speed $v$. Figure 6 shows the wave functions obtained after completing one full circle. In Fig. 6(a) corresponding to $v_{1}=150\,\mathrm{\mu m/s}$ (time to complete the circle is $t=0.314\,\mathrm{s}$), it is apparent that the vortex of the first component has retained its structure, and the density of the second component has not been strongly distorted. The value of $E$ has increased by only $3\%$ of the energy (per particle) gap between the ground state and the first excited one. Here the velocity is chosen to be lower than the speed of sound in the second component, which equals to $v_{\rm s}=\hbar/(m\xi_{2}\sqrt{2})=400\,\mathrm{\mu m/s}$. Moving the vortex at twice the speed, $v_{2}=2v_{1}$, results in setting the whole of the second component in circular motion. The distortion of the density is apparent in Fig. 6(b).
@@
Finally, the case of supersonic movement is studied in Fig. 6(c) corresponding to the movement speed $v_{3}=2v_{\rm s}$. The density profiles of both components are strongly distorted and the energy is increased 1.5 times. Nevertheless, the first component retains its signature phase profile, and no additional vortices are created in the second component in the process. Let us stress that in all three cases, the third component remained essentially unpopulated, and the total norm of the wave functions was conserved on the level of the numerical accuracy used, meaning that, in practice, no loss of atoms took place.
It is also instructive to use the localized vortex in the first component as a probe for the critical superfluid velocity $v_{\rm c}$ of the second component. In this case, the vortex acts as an impurity that can move without resistance provided its speed remains below $v_{\rm c}$. According to Landau’s criterion, the phononic excitation spectrum of the condensate sets the critical velocity equal to the sound velocity [41]. However, both experimental results [42, 43, 44, 45] and detailed theoretical studies [46, 47, 48] indicate that $v_{\rm c}$ is typically lower, depending on factors such as impurity size [44], the nature of the impurity potential (attractive or repulsive) [46], and the impurity trajectory (linear or circular) [46, 48]. To measure $v_{\rm c}$ in our system, we repeat the circular trajectory simulations, moving the beams for a fixed duration of $0.1\,\mathrm{s}$. We monitor the total energy per particle, $E(t)$, and fit it with a linear function to obtain the heating rate $\kappa=\mathrm{d}E/\mathrm{d}t$. Since the first component’s population is much smaller than that of the second, $E(t)$ primarily reflects the energy of the atoms in the second component. The results for velocities from $0.1v_{\rm s}$ to $1.2v_{\rm s}$ are shown in Fig. 7, revealing a clear threshold behavior where $\kappa$ increases rapidly once the velocity exceeds a certain value.
We quantified this threshold by fitting the data to
$\kappa(v)=\kappa_{0}+b\max(v^{2}-v_{{\rm c}}^{2},0),$ (20)
where $\kappa_{0}$, $b$, and $v_{\rm c}$ are fit parameters. This form is suitable for 2D systems [49] and has proven effective in experimental analyses [44, 47]. The least-squares fit shown in Fig. 7 yields $v_{\rm c}=0.8v_{\rm s}$, which is lower than the sound velocity and consistent with the above observations. The precise dissipation mechanisms warrant further study and will be addressed in a future work.
Figure 7: Condensate heating rate as a function of the movement speed of the beams.
@@ V Conclusions
In summary, we have studied the interaction of a two-component BEC mixture with the light fields in a $\Lambda$-type configuration and investigated the stationary states of the dark-state manifold. The angular momentum of $\nu\hbar$ per photon carried by one of the two fields leads to either one or both components being in a vortex state, with their vorticities differing by $\nu$ units. In the regime $a\ll R$, the structure of the stationary states is dictated by the vector potential term. In this limit, $\boldsymbol{A}$ tends to $\frac{\nu}{\rho}\boldsymbol{e}_{\varphi}$ and its contribution to the velocity field (and hence to the kinetic energy) can be canceled if the dark state is described by a wave function $f(\rho)\mathrm{e}^{-\mathrm{i}\nu\varphi}$. We have demonstrated that the lowest-energy state of the dark-state manifold indeed has this form. In this case, the first component (the one interacting with the LG beam) contains a vortex of vorticity $\nu$, while the second component is vortex-free. The core of the former vortex coincides with the center of the beams, while the density profile of the vortex demonstrates a strong degree of localization as the density falls off as $1+(\rho/a)^{2}]^{-1/2}$ away from the vortex core. Such a vortex can be moved around by moving the laser beams. Provided the movement speed is less than approximately half the speed of sound in the second component, the shape of the vortex retains its structure during the movement, and the density of the second component does not get distorted. More specifically, moving the vortex at a speed below the critical value of about $0.8v_{\rm s}$ causes almost no heating of the condensate, indicating its superfluid nature.
@@ Acknowledgements.
Y. B. would like to acknowledge the outstanding hospitality received during his internship at the Okinawa Institute of Science and Technology Graduate University, Quantum Systems Unit, where work on this project was initiated. This project was supported by the Research Council of Lithuania (RCL) grant No. S-LJB-24-2 and the JSPS Bilateral Program No. JPJSBP120244202. This work has also received funding from COST Action POLYTOPO CA23134, supported by COST (European Cooperation in Science and Technology).
@@
Our calculations were carried out in the dark-state regime, where the amplitudes of $\Phi_{3}$ were vanishingly small. Consequently, the non-Hermitian nature of the dynamics arising from the imaginary decay term $\Gamma$ was effectively negligible, allowing us to adopt the conventional normalization used for Hermitian systems, $\sum_{i=1}^{3}\int\mathrm{d}\boldsymbol{r}\,|\Phi_{i}|^{2}=1$ (the sum can equally be taken over $\Phi_{1}$ and $\Phi_{2}$ only). Furthermore, the energies and chemical potentials of the stationary states had vanishingly small imaginary parts, implying that the wave function norm remained essentially constant over time. We note that in our numerical calculations, the wave functions were normalized to unity rather than to the total particle number.
@@
$\begin{split}\boldsymbol{A}&=-\frac{\nu}{a}\frac{\left(\frac{\rho}{a}\right)^{2\nu-1}}{1+\left(\frac{\rho}{a}\right)^{2\nu}}\boldsymbol{e}_{\varphi},\\ U&=2\left(\frac{\nu}{a}\right)^{2}\frac{\left(\frac{\rho}{a}\right)^{2\nu-2}}{[1+\left(\frac{\rho}{a}\right)^{2\nu}]^{2}}.\end{split}$ (A4)